Miltmar Corporation will pay a year-end dividend of $4.00 and dividends thereafter are expected to grow at a constant rate of 3% per year. The risk-free rate is 5% and the expected return on the market portfolio is 14%. The stock has a beta of 0.65. A) Calculate the market capitalization rate. B) What is the intrinsic value of its stock?
Answer #1

a). Market Capitalization Rate(r) = Risk-free Rate + [Beta * (Expected Market Return - Risk-free Rate)]

= 5% + [0.65 * (14% - 5%)] = 5% + [0.65 * 9%] = 5% + 5.85% = 10.85%

b). Intrinsic Value = [Current Dividend * (1 + g)] / [r - g]

= [$4 * (1 + 0.03)] / [0.1085 - 0.03] = $4.12 / 0.0785 = $52.48